A rectangular garden measures 12m by 5m, a path of uniform width runs along one side and one end.
If the total area of the garden and path is 98 m². Find the width of the path

Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)

A rectangular garden measures 12m by 5m, a path of uniform width runs along one side and one end.
If the total area of the garden and path is 98 m². Find the width of the path
Let uniform width of path be x
Add x to length and breadth
the dimensions will be (12+x) and (5+x)
Area of plot = 98 m^2
(x+12)(x+5) =98
x^2+17x+60=98
x^2+17x-38 =0
(x+19)(x-2)=0
x= -19 or x=2

Ignore negative
width = 2 m
.







Answer by greenestamps(13198)



A quick informal solution, if formal algebra is not required....

The difference between the length and width of the garden is 12-5 = 7m; since the path of uniform width is on one side and one end, the difference between the length and width of garden plus path is still 7m.

Since the area of garden and path is 98 m^2, quick mental arithmetic says the dimensions of garden plus path are 14x7; that makes the width of the path 2m.